If the currents in the three parallel branches of an ac circuit are given as I1 = 25∠30°A, I2 = 25e-jΠ/6A and I3 = (50 + j 50)A, express the total current in the form

(i) I ∠Φ and

(ii) Im sin (ωt + Φ).


1
Expert's answer
2021-02-28T07:37:27-0500

"I_1=25\angle30\degree\space A=25(cos\dfrac{\pi}{6}+jsin\dfrac{\pi}{6})=(21.65+j12.5)\space A"

"I_2=25e^{-j\pi/6}\space A=25(cos\dfrac{-\pi}{6}+jsin\dfrac{-\pi}{6})=(21.65-j12.5)\space A"

"I_3=50+j50\space A"


(a) "I=I_1+I_2+I_3"

"\Rightarrow I=(21.65+21.65+50)+j(12.5-12.5+50)"

"\Rightarrow I=93.3+j50"

Magnitude of "I=\sqrt{(93.3)^2+(50)^2}"

"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=105.85"

Position of "I" with respect to "x" axis is "\theta=tan^{-1}\dfrac{50}{93.3}"

"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=28.18\degree"

"I=105.85\angle28.18\degree\space A"


(b) "I_m=105.85\space A"

"\phi=28.18\degree"

Assuming frequency of AC supply to be 50 Hz

"\omega=2\pi f=2\pi\times50=314s^{-1}"

"I=I_m(\omega t+\phi)"

"I=105.85sin(314t+28.18\degree)\space A"